Which function when graphed would include the point(3,0.125)?
A. f(x)=x^0.5
B. f(x)=0.5x
C. f(x)=0.5/x
D. f(x)=0.5^x

Answer by DrBeeee(684)

The short answer is D.
A) f(3) = sqrt(3) is not 1/8
B) f(3) = 3/2 is not 1/8
C) f(3) = 1/(2*3) is not 1/8
D) f(3) = (1/2)^3 is 1/8